The perimeter of a rectangle equals the sum of the length of its four sides.if the width of a rectangle is 6 inches and its perimeter is 44 inches, find the length of the rectangle.
Answer by ReadingBoosters(3246)

p = 2l+2w = 2(l+w) = 44
w = 6
44 = 2(6+l)
44 = 12 + 2l
44-12 = 2l
32 = 2l
16 in = length
...................
Delighted to help
-Reading Boosters
Wanting for others what we want for ourselves.